# Neural Networks and (Virtual) Extended Formulations

### v3

@@ 4.1. Neural Networks are Virtual Extended Formulations
Let $\mathcal{N}_{0}$ be the maxout network of size $s$ representing $f$. Without loss of generality, we assume that the underlying graph of $\mathcal{N}_{0}$ is transitively closed. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
@@
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first p1p-1 neurons unchanged. Now consider the weights $w_{uv}^{i}$ of $v\coloneqq v_{p}$. We split them into a positive and negative part, that is, $w_{uv}^{i}=a_{uv}^{i}-b_{uv}^{i}$ with $a_{uv}^{i}=\max\{w_{uv}^{i},0\}$, $b_{uv}^{i}=\max\{-w_{uv}^{i},0\}$. Observe that
@@
We continue that way until we obtain $\mathcal{N}_{s-1}$, where all weights except those of the output neuron are nonnegative. Performing the same splitting of the weights into positive and negative parts as before and looking at the expression of (3) for the output neuron $v\coloneqq v_{s}$, we define $z_{v}\eqqcolon g-h$ according to (3). While $g$ is already a maxout expression with only nonnegative weights, observe that $h$ can be artificially converted into such a maxout expression by taking the maximum over the same linear expression in the $z_{u}$-terms $k$ times.
@@ 4.2. Optimizing over Virtual Extended Formulations
To this end, assume that we have polytopes $P+Q=R$ with small extended formulations for $Q$ and $R$. Switching to support functions implies that for all objective directions $c\in\mathbb{R}^{d}$ we obtain $\max_{x\in P}c^{\top}x+\max_{x\in Q}c^{\top}x=\max_{x\in R}c^{\top}x$. Thus, to obtain the optimal objective value, it is sufficient to have small extended formulations for $Q$ and $R$, as we can really just optimize over $Q$ and $R$ and subtract the results. We now argue that this is basically also true for the solution itself.
@@ 4.3. Extension Complexity and Minkowski Sums
For a better understanding of virtual extension complexity, the behavior of extension complexity under Minkowski sum is crucial. It is well-known that $xc(R)\leq xc(P)+xc(Q)$ for $P+Q=R$. This follows from plugging the extended formulations for $P$ and $Q$ into the definition of the Minkowski sum. However, for bounding virtual extension complexity, one would need to bound $\operatorname{xc}(P)$ in terms of $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$. In the following, we present a class of examples demonstrating that one cannot simply focus on $\operatorname{xc}(R)$ for this, but also needs to take $\operatorname{xc}(Q)$ into account.
@@
Considering a matroid polytope $P$ with high extension complexity from the latter proof such that $P+Q=\Pi_{n}$ for some other generalized permutahedron $Q$, we cannot exclude that $Q$ has also high extension complexity. Therefore, without a better understanding of $Q$, this does not yield any useful bound on $\operatorname{vxc}(P)$.
On a final note, we would like to emphasize that in the definition of the virtual extension complexity, it is important that $Q$ and $R$ are polytopes and not potentially unbounded polyhedra. A simple example is given by the choice of $Q=R=\mathbb{R}^{d}$, which would immediately imply $\operatorname{vxc}(P)=0$ for any polyhedron $P$, if it was allowed. The important algebraic property that we need to make sure that we can recover $P$ from $Q$ and $R$ is the cancellation property: For polytopes $A,B,C$, it is true that $A+C=B+C$ implies $A=B$. The same is not true for potentially unbounded polyhedra.
@@ 5. Conclusions
Closely related, and from the perspective of polyhedral theory and combinatorial optimization, the most intriguing open question is: Can $\operatorname{vxc}(P)$ be (much) smaller than $\operatorname{xc}(P)$? If the answer is no, then this would imply strong lower bounds on neural networks. On the other hand, if the answer is yes, it implies that solving two linear programs and taking the difference is (much) more powerful than just solving one linear program. In fact, we are not even aware of an example for which we can prove that $\operatorname{vxc}(P)<\operatorname{xc}(P)$. On the other hand, as for the usual extension complexity, we do not expect $\operatorname{vxc}(P)$ to be polynomial for any polytope $P$ associated with an NP-hard optimization problem. An obvious, but seemingly challenging approach to lower-bound virtual extension complexity is to take any previously successful method for lower-bounding extension complexity and try to adapt it such that it also works in the virtual case.
@@ Acknowledgments
Part of this work was completed while Christoph Hertrich was affiliated with Université Libre de Bruxelles, Belgium, and received support by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No 101153187—NeurExCo.

### v4

@@ 4.1. Neural Networks are Virtual Extended Formulations
Let $\mathcal{N}_{0}$ be the maxout network of size $s$ representing $f$. Without loss of generality, we assume that the underlying graph of $\mathcal{N}_{0}$ is transitively closed. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network according to our definition of size [^6] As above, this might however increase the number of arcs and parameters of the network.. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
@@
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we leave the first p1p-1 neurons unchanged. Now consider the weights $w_{uv}^{i}$ of $v\coloneqq v_{p}$. We split them into a positive and negative part, that is, $w_{uv}^{i}=a_{uv}^{i}-b_{uv}^{i}$ with $a_{uv}^{i}=\max\{w_{uv}^{i},0\}$, $b_{uv}^{i}=\max\{-w_{uv}^{i},0\}$. Observe that
@@
We continue that way until we obtain $\mathcal{N}_{s-1}$, where all weights except those of the output neuron are nonnegative. For the output neuron, we perform the same splitting operation of the weights into positive and negative parts as for all previous neurons, obtaining (3) for $v\coloneqq v_{s}$. This allows us to define $g$ and $h$ via the right-hand side of (3): $g$ is the maximum expression and $h$ is the part behind the minus sign. By definition, this implies $f=z_{v_{s}}=g-h$. Now observe that by our construction, $g$ is already a maxout expression with only nonnegative weights. Furthermore, observe that $h$ can be artificially converted into such a maxout expression by taking the maximum over the same linear expression in the $z_{u}$-terms $k$ times.
@@ 4.2. Optimizing over Virtual Extended Formulations
To this end, assume that we have polytopes $P+Q=R$ with small extended formulations for $Q$ and $R$. Switching to support functions implies that for all objective directions $c\in\mathbb{R}^{d}$ we obtain $\max_{x\in P}c^{\top}x+\max_{x\in Q}c^{\top}x=\max_{x\in R}c^{\top}x$; see also Figures 1 and 4. Thus, to obtain the optimal objective value, it is sufficient to have small extended formulations for $Q$ and $R$, as we can really just optimize over $Q$ and $R$ and subtract the results. We now argue that this is basically also true for the solution itself.
@@ 4.3. Extension Complexity and Minkowski Sums
For a better understanding of virtual extension complexity, the behavior of extension complexity under Minkowski sum is crucial. It is well-known that $\operatorname{xc}(R)\leq\operatorname{xc}(P)+\operatorname{xc}(Q)$ for $P+Q=R$. This follows from plugging the extended formulations for $P$ and $Q$ into the definition of the Minkowski sum. However, for bounding virtual extension complexity, one would need to bound $\operatorname{xc}(P)$ in terms of $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$. In the following, we present a class of examples demonstrating that one cannot simply focus on $\operatorname{xc}(R)$ for this, but also needs to take $\operatorname{xc}(Q)$ into account.
@@
Considering a matroid polytope $P$ with high extension complexity from the latter proof such that $P+Q=\Pi_{n}$ for some other generalized permutahedron $Q$, we cannot exclude that $Q$ also has high extension complexity. Therefore, without a better understanding of $Q$, this does not yield any useful bound on $\operatorname{vxc}(P)$.
On a final note, we would like to emphasize that in the definition of the virtual extension complexity, it is important that $Q$ and $R$ are polytopes and not potentially unbounded polyhedra. A simple example is given by the choice of $Q=R=\mathbb{R}^{d}$, which would immediately imply $\operatorname{vxc}(P)=0$ for any polyhedron $P$, if this were allowed. The important algebraic property that we need to make sure that we can recover $P$ from $Q$ and $R$ is the cancellation property: For polytopes $A,B,C$, it is true that $A+C=B+C$ implies $A=B$. The same is not true for potentially unbounded polyhedra.
@@ 5. Conclusions
Closely related, and from the perspective of polyhedral theory and combinatorial optimization, the most intriguing open question is the following.
Question 5.1.
Can $\operatorname{vxc}(P)$ be (much) smaller than $\operatorname{xc}(P)$?
We do not see any clear evidence in either direction for the resolution of 5.1. If the answer is no, then this would imply strong lower bounds on neural networks. For that, it would even be sufficient to prove that $\operatorname{xc}(P)$ cannot be exponentially larger than $\operatorname{vxc}(P)$. On the other hand, if the answer to 5.1 is yes, it would imply that solving two linear programs and taking the difference is (much) more powerful than just solving one linear program.
In fact, we are not even aware of an example for which we can prove that $\operatorname{vxc}(P)<\operatorname{xc}(P)$. A first step into this direction is the result by Hertrich et al. (2026), showing that for $P$ being the base polytope of a regular matroid on $n$ elements, $\operatorname{vxc}(P)\in O(n^{3})$, while the best known upper bound on $\operatorname{xc}(P)$ is $O(n^{6})$ (Aprile and Fiorini, 2022). However, it is conceivable that the $\operatorname{xc}$ upper bound is suboptimal and there might be no actual separation.
As for the usual extension complexity, we do not expect $\operatorname{vxc}(P)$ to be polynomial for any polytope $P$ associated with an NP-hard optimization problem. An obvious, but seemingly challenging approach to lower-bound virtual extension complexity is to take any previously successful method for lower-bounding extension complexity and try to adapt it such that it also works in the virtual case. Most such approaches in the literature rely on the theorem by Yannakakis (1991) stating that $\operatorname{xc}(P)$ equals the nonnegative rank of the slack matrix. Therefore, a first step would be to establish a similar, even if weaker, characterization of $\operatorname{vxc}(P)$ in terms of the slack matrix. Beyond this, we think the most promising route might be to focus on lower bounds based on counting arguments, similar to Rothvoß (2013), with the aim to obtain an analogue of Theorem 3.7 for $\operatorname{vxc}$.
@@ Acknowledgments
Part of this work was completed while Christoph Hertrich was affiliated with Université Libre de Bruxelles, Belgium, and received support from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No 101153187—NeurExCo.